OpenStudy (anonymous):

Write is 75.6% as a fraction in simplest form?

OpenStudy (anonymous):

75 3 _ 5 I think, hope It's right.

OpenStudy (phi):

you can "get rid" of the % sign if you drop the % and instead divide by 100 like this \[ 75.6\% \rightarrow \frac{75.6}{100} \] you do not want a decimal in your fraction so multiply top and bottom by 10 \[ \frac{75.6\cdot 10}{100\cdot 10} \] simplify that to \[ \frac{756}{1000} \] we are not yet done, because we can simplify some more. can you finish ?

OpenStudy (anonymous):

700

OpenStudy (phi):

@tatumrocks I would divide top and bottom by 4 to reduce the fraction to simplest form.